OpenStudy (anonymous):

evaluate the logarithmic equation log6 1296

OpenStudy (anonymous):

6^x=1296

OpenStudy (lgbasallote):

\[\huge \log _6 (1296) \implies \log (1296) \div \log (6)\] you can use your calculator to solve

OpenStudy (anonymous):

@lgbasallote how do u type like that

OpenStudy (lgbasallote):

benefits of level 99s

OpenStudy (anonymous):

So, only 99s can write those